# Model-Based Transfer Learning for Contextual Reinforcement Learning

### v2

@@ 5.3 Continuous control benchmark experiments
Table 2: Comparative performance of different methods on standard control CMDPs
Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
@@
Pendulum Length 0.7383 0.6830 0.7270 0.7092 0.7327 0.7697 0.7969
Pendulum Mass 0.6237 0.5793 0.6329 0.6092 0.6408 0.6827 0.7132
Pendulum Timestep 0.8135 0.7247 0.7989 0.7541 0.8177 0.8141 0.8801
Cartpole Mass of Cart 0.9466 0.7153 0.7221 0.7516 0.6501 0.8212 0.9838
Cartpole Length of Pole 0.9110 0.5441 0.8121 0.8428 0.8217 0.9124 0.9875
Cartpole Mass of Pole 0.9560 0.6073 0.8858 0.7909 0.8744 0.9351 1.0000
BipedalWalker Gravity 0.9281 0.7898 0.9330 0.9494 0.9359 0.9393 0.9674
BipedalWalker Friction 0.9317 0.9051 0.965 0.9664 0.9645 0.9713 0.9778
BipedalWalker Scale 0.8694 0.7452 0.8605 0.8496 0.8792 0.8886 0.9107
HalfCheetah Gravity 0.6679 0.6292 0.8542 0.8634 0.8663 0.9073 0.9544
HalfCheetah Friction 0.6693 0.7242 0.8567 0.8703 0.8591 0.9274 0.9663
HalfCheetah Stiffness 0.6561 0.7007 0.8533 0.7817 0.8785 0.9146 0.9674
Average 0.8093 0.6957 0.8251 0.8116 0.8267 0.8736 0.9255
@@
Figure 5: Continuous control CMDP results. Method comparison of normalized performance over $N$ tasks. The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 50x fewer samples needed.
To probe the generality of MBTL, we utilize context-extended versions of standard RL environments from CARL benchmark library [6] to evaluate our methods under varied contexts. For the Cartpole domain, we considered CMDPs with varying cart mass, pole length, and pole mass. In Pendulum, we vary the timestep duration, pendulum length, and pendulum mass. The BipedalWalker was tested under varying friction, gravity, and scale. In HalfCheetah domain, we manipulated friction, gravity, and stiffness parameters. These variations critically influence the dynamics and physics of the environments. The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times ($N=100$), enabling a comprehensive analysis of transfer learning under drastically different conditions. We provide more details in Appendix A.4.
@@
Results. The results summarized in Table 2 demonstrate MBTL’s sample efficiency and competitive performance across diverse control domains, often closely trailing the Oracle only with a small number of tasks. With the exception of the Cartpole domain, MBTL generally shows superior performance. Figure 5 shows that MBTL demonstrates rapid learning curves, reaching high normalized performance with relatively few samples. Specifically, Figure 6 illustrates the detailed process of how MBTL utilizes GP for performance estimation and chooses the next source task that maximizes the acquisition function that evaluates the expected improvement of generalized performance. MBTL achieves comparable performance to multi-task or independent baselines with 8-50x fewer samples, highlighting its effectiveness in reducing training requirements.
@@ 6 Related work
Contextual Reinforcement Learning. Robustness and generalization challenges in DRL are generally addressed by a few common techniques in the literature. The broader umbrella of such methods falls under CRL [6]. CRL utilizes side information about the problem variations to improve the generalization and robustness. In particular, CRL formalizes generalization in DRL using CMDPs [15, 32, 6], which incorporate context-dependent dynamics, rewards, and initial state distributions into the formalism of MDPs. This provides a structured framework to study generalization in DRL [15, 32]. The contexts of CMDPs are not always visible during training [22]. When they are visible, they can be directly used as side information by conditioning the policy on them [38].
Multi-task learning. A common approach for solving CRL is multi-task learning, which aims to leverage common structure between tasks. The structure between tasks has been exploited using policy sketches for task structuring [2] and by sharing a distilled policy that captures common behavior across tasks [44]. One of the known challenges in multi-task learning is that when the context features are not visible, the resulting CMDP becomes a partially observable CMDP [22, 10] and can be challenging for multi-task learning. Moreover, another challenge is negative transfer, where training from tasks that are too different results in training instability or failure [21, 41, 42]. In this work, we use multi-task learning as a baseline in evaluating our methods.
Zero-shot transfer and policy reuse. Zero-shot transfer is another commonly used technique that focuses on adapting models trained in one environment to perform effectively in unseen settings without further training [22]. Li et al. [25] introduces context-aware policy reuse, which adapts multiple source policies based on context, enhancing generalization in complex environments. Rezaei-Shoshtari et al. [34] utilize hypernetworks to approximate an RL algorithm as a mapping from parameterized CMDP to a family of near-optimal solutions to achieve zero-shot transfer across varying task conditions. Sinapov et al. [37] use meta-data to learn inter-task transferability to learn the expected benefit of transfer given a source-target task pair. Bao et al. [3] propose a metric for evaluating transferability based on information-theoretic feature representations across tasks. Another popular class of techniques for robust and generalizable DRL is domain adaptation strategies. It is often used in making learned policies adapt to new domains with varying dynamics, observations, or rewards [5]. For instance, Xing et al. [46] first learn a latent unified state representation for different domains and subsequently train DRL agents in the source domain based on that. However, our focus in this work centered around in-distribution generalization, whereas domain adaptation techniques are commonly used for out-of-distribution generalization.
@@
Source task selection. In the context of transfer learning, selecting appropriate source tasks is crucial. Li and Zhang [24] proposes an optimal online method for dynamically selecting the most relevant single source policy in reinforcement learning. Beyond RL, Meiseles and Rokach [28] emphasizes structural alignment in time-series source models to prevent performance degradation, while Poth et al. [33] finds that selecting aligned intermediate tasks in natural language processing boosts transfer effectiveness. Together, these studies show the importance of context-aligned source tasks selection to improve transferability and reduce negative transfer.
@@ 7 Conclusion
This study introduces a method called Model-based Transfer Learning (MBTL), which layers on top of existing RL methods to effectively solve CMDPs. Rather than independent or multi-task training, which trains $N$ or 1 models, respectively, MBTL intelligently selects an intermediate number of models to train. MBTL has two key components: an explicit model of the generalization gap and a Gaussian process component to estimate training performance. MBTL achieves up to 50x improved sample efficiency on standard and real-world benchmarks. Furthermore, MBTL achieves sublinear regret in the number of training tasks. A limitation is that MBTL is designed for a single-dimensional context variation. Promising directions of future work include studying high-dimensional context spaces and out-of-distribution generalization, as well as the development of new real-world CMDP benchmarks.
@@ A.1 Notation
Figure 9 helps understand the discrepancy between the observed generalized performance and the predicted one. Figure 10 illustrates how to calculate the marginal improvement of expected generalized performance ($\hat{V}(x;x_{1:k-1})-V(x;x_{1:k-1})$).
@@ A.4 Experiment details
A.4.1 Details about Gaussian process (GP)
In our study, we conducted hyperparameter tuning experiments for Gaussian Process (GP) regression to optimize its performance. Specifically, we varied the noise standard deviation over the set $\{0.001,0.01,0.1,1\}$ and the number of restarts for the optimizer over the set $\{5,6,7,8,9,10,11,12,13,14,15\}$ with several kernel configurations. We set noise standard deviation as $0.1$, the number of restarts for the optimizer over the set as $9$, and the kernel as the combination of a constant kernel (C) with the bound from 0.001 to 1000 and a radial basis function kernel (RBF) with a length scale ranging from 0.01 to 100. The noise level was set to the square of the noise standard deviation ($0.1^{2}$).
A.4.2 Results of table with standard deviation
Table 4: Comparative performance of different methods on context-variant traffic and benchmark CMDPs
Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
@@ A.4.2 Results of table with standard deviation
A.4.3 Detailed sample complexity comparison results
Table 5 presents a comparison of sample complexity required for MBTL to perform as good as the best generalization performance of baselines (Independent training and multi-task training) across various tasks in the CMDP. Each row lists a different domain, the specific context variation applied (e.g., changes in physical properties or environmental parameters), and two key values: $k^{*}$ and $N$, where $k^{*}$ represents the number of models required by MBTL to reach a performance level comparable to the baseline. This value is shown as a range (e.g., $[3,5,3]$), indicating results from three random seeds. $N$ represents the total number of contexts. The value $\frac{N}{k^{*}}$ helps illustrate the sample efficiency of MBTL.
@@ A.4.3 Detailed sample complexity comparison results
A.4.4 Details about traffic signal control
@@
Figure 11 showcases the layout of traffic networks used in a traffic signal control task with several lanes and a signalized intersection in the middle. The state space represents the presence of vehicles in discretized lane cells along the incoming roads. Actions determine which lane gets the green phase of the traffic signal, and rewards are based on changes in cumulative stopped time, the period when speed is zero. The global objective is to minimize the average waiting times at the intersection. Different configurations of intersections (e.g., road length, inflow, speed limits) are modeled to represent varying real-world conditions.
@@
We used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [35]. All experiments are done on a distributed computing cluster equipped with 48 Intel Xeon Platinum 8260 CPUs.
@@ Results
  • Speed Limit: Here, performance levels are relatively stable across all strategies except for the multitask training.
@@
A.4.5 Details about eco-driving control
@@
We also used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [35].
@@ Results
A.4.6 Details about advisory autonomy task
@@
Problem Definition: In a single-lane scenario, the state space includes the speeds of the ego and leading vehicles, along with the headway. For highway ramp scenarios, additional states cover the relative positions and speeds of adjacent vehicles. Actions vary by guidance type: for acceleration guidance, the action space is continuous, ranging from $-1$ to $1$; for speed guidance, it has ten discrete actions compared to the speed limit. Rewards are based on system throughput or average speed of all vehicles in the system.
Context Variations: We explore different durations of coarse-grained guidance holds to test various levels of human compatibility, adjusting the model based on observed driver behaviors and system performance.
@@ Results
A.4.7 Details of control tasks
@@
A.4.8 Details about Cartpole

### v3

@@ 5.3 Continuous control benchmark experiments
Table 2: Comparative performance of different methods on standard control CMDPs ($K=15$)
Benchmark (CMDP) Baselines Multi-policy Baselines MBTL Oracle
@@
Pendulum Length 0.7383 0.6830 0.7607 0.7601 0.7774 0.7755 0.7969
Pendulum Mass 0.6237 0.5793 0.6647 0.6794 0.6887 0.6667 0.7132
Pendulum Timestep 0.8135 0.7247 0.8331 0.8292 0.8497 0.8333 0.8801
Cartpole Mass of Cart 0.9466 0.7153 0.8961 0.9044 0.8299 0.9356 0.9838
Cartpole Length of Pole 0.9110 0.5441 0.9497 0.9484 0.9424 0.9488 0.9875
Cartpole Mass of Pole 0.9560 0.6073 0.9870 0.9927 0.9916 0.9813 1.0000
BipedalWalker Gravity 0.9281 0.7898 0.9654 0.9666 0.9656 0.9655 0.9674
BipedalWalker Friction 0.9317 0.9051 0.9739 0.9738 0.9738 0.9725 0.9778
BipedalWalker Scale 0.8694 0.7452 0.8910 0.8975 0.8990 0.8962 0.9107
HalfCheetah Gravity 0.6679 0.6292 0.9086 0.9000 0.9089 0.9214 0.9544
HalfCheetah Friction 0.6693 0.7242 0.9314 0.9457 0.9184 0.9225 0.9663
HalfCheetah Stiffness 0.6561 0.7007 0.9191 0.9097 0.9295 0.9205 0.9674
Average 0.8093 0.6957 0.8901 0.8923 0.8896 0.8950 0.9255
@@
Figure 6: Continuous control CMDP results. Method comparison of normalized performance over $N$ tasks. The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 43x improved sample efficiency.
To probe the generality of MBTL, we utilize context-extended versions of standard RL environments from CARL benchmark library [5] to evaluate our methods under varied contexts. For the Cartpole domain, we considered CMDPs with varying cart mass, pole length, and pole mass. In Pendulum, we vary the timestep duration, pendulum length, and pendulum mass. The BipedalWalker was tested under varying friction, gravity, and scale. In HalfCheetah domain, we manipulated friction, gravity, and stiffness parameters. These variations critically influence the dynamics and physics of the environments. The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times ($N=100$), enabling a comprehensive analysis of transfer learning under drastically different conditions. We provide more experimental details in Appendix A.4.
@@
Results. The results summarized in Table 2 demonstrate sample efficiency and competitive performance of multi-policy training including MBTL across diverse control domains, often closely trailing the Oracle only with a small number of trained policies. Figure 6 shows that with the exception of a few context variations, MBTL generally shows superior performance. Specifically, Figure 7 illustrates the detailed process of how MBTL utilizes GP for performance estimation and chooses the next source task that maximizes the acquisition function that evaluates the expected improvement of generalized performance. MBTL achieves comparable performance to multi-task or independent baselines with 5-43x fewer samples, highlighting its effectiveness in reducing training requirements.
@@ 6 Related work
Contextual Reinforcement Learning. Robustness and generalization challenges in DRL are generally addressed by a few common techniques in the literature. The broader umbrella of such methods falls under CRL, which utilizes side information about the problem variations to improve the generalization and robustness. In particular, CRL formalizes generalization in DRL using CMDPs [14, 32, 5], which incorporate context-dependent dynamics, rewards, and initial state distributions into the formalism of MDPs. The contexts of CMDPs are not always visible during training [23]. When they are visible, they can be directly used as side information by conditioning the policy on them [39]. In this paper, we focus on a scenario where the learner can choose which context-MDP to train on. This contrasts with other CRL works that assume context-MDPs arrive from a fixed distribution.
Multi-task training. Multi-task methods can help address CRL by exploiting shared structure across tasks. Prior work has leveraged techniques such as policy sketches for task decomposition [2] and distilled policies that capture common behaviors [46]. However, a key limitation arises when the context is unobserved, effectively transforming the CMDP into a partially observable setting [23, 9], which complicates multi-task training. Another challenge is negative transfer, wherein training on tasks that are too dissimilar leads to instability or outright failure [22, 42, 44]. Although more recent multi-task approaches such as MOORE [16] and PaCo [43] have shown promise, they often focus on discrete task sets and are thus less suited to CRL, where tasks span a broad continuum of contexts. In this work, we include multi-task learning as a baseline to benchmark our methods.
Zero-shot transfer and policy reuse. Zero-shot transfer—where models trained for one environment directly perform in new, unseen settings without additional training [23]—is an important strategy in CRL settings. For solving CMDP problems, prior works attempted to utilize zero-shot transfer to solve CMDP problems by approximation on RL algorithm and hypernetworks that maps from parameterized CMDP to a family of near-optimal solutions [35]. Sinapov et al. [38] use meta-data to learn inter-task transferability to learn the expected benefit of transfer given a source-target task pair. Bao et al. [3] propose a metric for evaluating transferability based on information-theoretic feature representations across tasks. Taken together, these approaches highlight the importance of policy reuse, where efficiently selecting or adjusting a pre-trained policy accelerates learning and improves robustness in new contexts.
@@
Source task selection. In the context of transfer learning, selecting appropriate source tasks is crucial. Li and Zhang [25] proposes an optimal online method for dynamically selecting the most relevant single source policy in reinforcement learning. Beyond RL, Meiseles and Rokach [28] emphasizes structural alignment in time-series source models to prevent performance degradation, while Poth et al. [33] finds that selecting aligned intermediate tasks in natural language processing boosts transfer effectiveness. Building upon these insights, we formulate the source task selection problem for CRL, enabling zero-shot transfer by estimating training performance online and leveraging structural generalization across context variations.
@@ 7 Conclusion
This study introduces a method called Model-based Transfer Learning (MBTL), which layers on top of existing RL methods to effectively solve CMDPs. Rather than independent or multi-task training, which trains $N$ or 1 models, respectively, MBTL intelligently selects an intermediate number of models to train. MBTL has two key components: an explicit model of the generalization gap and a Gaussian process component to estimate training performance. MBTL achieves up to 43x improved sample efficiency on standard and real-world benchmarks. Furthermore, MBTL achieves sublinear regret in the number of training tasks. A limitation is that MBTL is designed for a single-dimensional context variation with a reliance on the explicit similarity of context variables. Promising directions of future work include studying high-dimensional context spaces and formalizing task similarity, as well as the development of new real-world CMDP benchmarks.
@@ A.1 Notation
Figure 10 helps understand the discrepancy between the observed generalized performance and the predicted one. Figure 11 illustrates how to calculate the marginal improvement of expected generalized performance ($\hat{V}(x;\pi_{1:k-1})-V(x_{1:k-1})$).
@@ A.4 Experiment details
A.4.1 Details about Gaussian process (GP) Regresstion
We use the GaussianProcessRegressor implementation from scikit-learn, which follows Algorithm 2.1 of [48]. Specifically, we construct a kernel by multiplying a constant kernel
$C(\theta)=1.0,\quad\theta\in(10^{-3},10^{3}),$
by a radial basis function (RBF) kernel
$k_{\mathrm{RBF}}(\mathbf{x},\mathbf{x}^{\prime};\ell)=\exp\Bigl{(}-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2\ell^{2}}\Bigr{)}$
with an initial length scale $\ell=1.0$ (constrained to lie in the range $[10^{-2},10^{2}]$). To determine the hyperparameters, we begin by generating synthetic data that aligns with our modeling assumptions, including constant training performance and a linear generalization gap, while introducing noise to degrade generalization performance by up to 10%, sampled from a uniform distribution. We vary the GP hyperparameters, including noise standard deviation over the set $\{0.001,0.01,0.1,1\}$, the number of restarts for the optimizer over $\{5,6,\dots,15\}$, and explore several kernel configurations on the synthetic data. We then select the hyperparameter configuration that maximizes the average predictive performance. Specifically, we choose a noise standard deviation of $\sigma=0.001$ and perform $15$ random restarts of the hyperparameter optimizer to reduce the risk of convergence to poor local minima. We use the same GP hyperparameter configuration across all experiments and benchmarks.
A.4.2 Accuracy of generalization gap assumption
In Figure 12, we report the Pearson correlation between the observed generalization gap and the estimated gap under our linear assumption (Assumption 1). Each histogram shows how strongly the two measures align across various tasks in standard control (blue) and traffic (red) benchmarks. Many tasks cluster around moderate positive correlations (0.3–0.5), suggesting that a linear function of context similarity can reasonably capture the gap in most scenarios. However, certain tasks—such as Eco-driving—exhibit higher correlations (above 0.6), whereas others—such as HalfCheetah—are closer to 0, indicating that the assumption holds more effectively in some domains than in others.
Figure 12: Accuracy of linear generalization gap assumption. Pearson correlation analysis on the observed generalization gap and the estimated gap under our linear assumption.
A.4.3 Results of table with standard deviation
Table 4: Comparative performance of different methods on context-variant traffic and benchmark CMDPs ($K=15$)
Benchmark (CMDP) Baselines Multi-policy Baselines MBTL Oracle
@@ A.4.3 Results of table with standard deviation
A.4.4 Detailed sample complexity comparison results
Table 5 presents a comparison of sample complexity required for MBTL to perform as good as the best generalization performance of baselines (independent training and multi-task training) across various tasks in the CMDP. Each row lists a different domain, the specific context variation applied (e.g., changes in physical properties or environmental parameters), and two key values: $k^{*}$ and $N$, where $k^{*}$ represents the number of models required by MBTL to reach a performance level comparable to the baseline. This value is shown as a range (e.g., $[3,5,3]$), indicating results from three random seeds. $N$ represents the total number of contexts. The value $\frac{N}{k^{*}}$ helps represent the sample efficiency of MBTL.
@@ A.4.4 Detailed sample complexity comparison results
A.4.5 Details about traffic signal control benchmark
@@
Figure 13 showcases the layout of traffic networks used in a traffic signal control task with several lanes and a signalized intersection in the middle. The state space represents the presence of vehicles in discretized lane cells along the incoming roads. Actions determine which lane gets the green phase of the traffic signal, and rewards are based on changes in cumulative stopped time, the period when speed is zero. The global objective is to minimize the average waiting times at the intersection. Different configurations of intersections (e.g., road length, inflow, speed limits) are modeled to represent varying real-world conditions. To represent various real-world conditions, we vary factors such as road length, inflow rate, and speed limits from 0.1 to 5 times; by default, the road length is 750 meters, the flow rate is 1000 vehicles per hour, and the speed limit is 13.89 m/s.
@@
We used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [36]. We utilized the default implementation of the PPO algorithm with default hyperparameters provided by the Stable-Baselines3 library [34]. All experiments are done on a distributed computing cluster equipped with 48 Intel Xeon Platinum 8260 CPUs.
License
Traffic signal control benchmark falls under MIT License.
Potential of multi-policy training and zero-shot transfer
Figure 14: Normalized performance of three DRL-based methods—Oracle Transfer (gray), independent training (orange), and multi-task training (green)—under different traffic-signal benchmarks: flow rate (left), speed limit (middle), and lane length (right). While independent and multi-task training approaches exhibit higher variance and reduced asymptotic performance, Oracle Transfer benefits from zero-shot transfer with full information, demonstrating more stable and generally higher performance.
Figure 14 shows how each approach adapts to variations in flow rate, speed limit, and lane length for a traffic signal control benchmark. The y-axis shows normalized performance, with higher values indicating better control policies. Oracle Transfer consistently achieves superior performance across these different settings, owing to its ability to leverage full task information in a zero-shot transfer manner. By contrast, independent and multi-task training exhibit more pronounced performance drops and greater instability when faced with shifts in problem parameters, underscoring the challenges of generalizing policies in traditional DRL approaches.
@@ Results
  • Speed Limit: Here, performance levels are relatively stable across all strategies except for the multi-task training.
@@
A.4.6 Details about eco-driving control benchmark
@@
We also used the microscopic traffic simulation called Simulation of Urban MObility (SUMO) [26] v.1.16.0 and PPO for RL algorithm [36]. For detailed experimental details and RL hyperparameter configurations, please refer to [20, 21].
License
Eco-driving benchmark falls under MIT License [21].
Potential of multi-policy training and zero-shot transfer
Figure 1 shows how each RL training paradigm adapts to variations in green phase time, penetration rate, and inflow rate in the eco-driving control benchmark. Oracle Transfer remains the strongest method across all configurations, benefiting from zero-shot transfer. independent training shows unstable performance across different task variations, performance, while multi-task training lags behind. Overall, the trends highlight the advantage of leveraging zero-shot transfer in traffic CMDPs.
@@ Results
A.4.7 Details about advisory autonomy benchmark
@@
Problem Definition
In a single-lane scenario, the state space includes the speeds of the ego and leading vehicles, along with the headway. For highway ramp scenarios, additional states cover the relative positions and speeds of adjacent vehicles. Actions vary by guidance type: for acceleration guidance, the action space is continuous, ranging from $-1$ to $1$; for speed guidance, it has ten discrete actions compared to the speed limit. Rewards are based on system throughput or average speed of all vehicles in the system.
Context Variations
We explore different durations of coarse-grained guidance holds to test various levels of human compatibility, adjusting the model based on observed driver behaviors and system performance.
License
Advisory autonomy benchmark falls under MIT License [40].
@@
Potential of multi-policy training and zero-shot transfer
Figure 21: Normalized performance of Oracle Transfer, independent training, and multi-task training under Advisory Autonomy benchmark with human compatibility task variations.
Figure 21 shows that ring-road networks tend to yield higher performance and smaller gaps compared to highway ramp scenarios. In addition, independent training exhibits greater performance drop and variability due to training instability. Oracle Transfer retains clear potential improvements over other baselines.
@@ Results
A.4.8 Details of control benchmarks
@@
We utilized the default implementation of the DQN, A2C, and PPO algorithm with default hyperparameters provided by the Stable-Baselines3 library [34].
A.4.9 Details about Cartpole benchmark
Potential of multi-policy training and zero-shot transfer
Cartpole may be considered a simpler benchmark than traffic benchmarks, yet independent and multi-task training methods still face notable difficulty when faced with context variations. Figure 24 shows that Oracle Transfer performs at the highest performance across problem variations, while independent training or multi-task training shows a larger variance in performance.
Figure 24: Normalized performance of Oracle Transfer, independent training, and multi-task training in Cartpole benchmarks.